# Simplicity bias, algorithmic probability, and the random logistic map

### 31 Dec 2023 v1

@@ Abstract.
Simplicity bias is an intriguing phenomenon prevalent in various input-output maps, characterized by a preference for simpler, more regular, or symmetric outputs. Notably, these maps typically feature high-probability outputs with simple patterns, whereas complex patterns are exponentially less probable. This bias has been extensively examined and attributed to principles derived from algorithmic information theory and algorithmic probability. In a significant advancement, it has been demonstrated that the renowned logistic map $x_{k+1}=\mu x_{k}(1-x_{k})$, a staple in dynamical systems theory, and other one-dimensional maps exhibit simplicity bias when conceptualized as input-output systems. Building upon this foundational work, our research delves into the manifestations of simplicity bias within the random logistic map, specifically focusing on scenarios involving additive noise. This investigation is driven by the overarching goal of formulating a comprehensive theory for the prediction and analysis of time series.
Our primary contributions are multifaceted. We discover that simplicity bias is observable in the random logistic map for specific ranges of $\mu$ and noise magnitudes. Additionally, we find that this bias persists even with the introduction of small measurement noise, though it diminishes as noise levels increase. Our studies also revisit the phenomenon of noise-induced chaos, particularly when $\mu=3.83$, revealing its characteristics through complexity-probability plots. Intriguingly, we employ the logistic map to underscore a paradoxical aspect of data analysis: more data adhering to a consistent trend can occasionally lead to reduced confidence in extrapolation predictions, challenging conventional wisdom.
@@ 1. Introduction
Dynamical systems appear in many fields including financial, engineering, meteorological, medical, and other areas. Because of this, developing methods to predict or understand the behavior and trajectories of dynamical systems is an important area of applied mathematics research. Historically, such predictions have been based on modeling the relevant system (e.g., with ODEs). More recently, however, machine learning advances have given tools for the prediction and analysis of dynamical systems. The machine learning approach instead relies on directly extracting patterns from data rather than detailed modeling of the underlying mechanics [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19].
@@
Numerous studies of simplicity bias have been made in input-output maps, in which a general inverse relationship between the complexity of outputs and their respective probabilities has been observed [31, 32]. This work is based on information theory and computation results and was derived from arguments inspired by AIT, and can be viewed as an attempt to apply algorithmic probability in real-world settings [31]. The simplicity bias upper bound [31] gives a way to bound probability $P(x)$ of observing output pattern $x$ based essentially only on estimating the Kolmogorov complexity of the pattern in $x$. Simplicity bias analysis has been applied in many contexts, including machine learning and deep neural networks [33, 34, 35, 36, 37, 38, 39]. Simplicity bias and algorithmic probability are closely related to Occam’s razor [40], a fundamental basis of scientific reasoning and model selection, that simpler theories/models should be preferred over more complex theories/models provided they explain the data equally well.
The goal of this paper is to bridge the gap between machine learning and algorithmic information theory using dynamical systems, particularly focusing on the concept of simplicity bias within random logistic maps. We aim to extend the understanding of simplicity bias, traditionally studied in deterministic settings, to encompass the inherently stochastic nature of real-world systems as modeled by random dynamical systems. Our approach is twofold: first, to elucidate the manifestation of simplicity bias in the random logistic map under various conditions and noise levels, thereby providing insights into the behavior of these systems; second, to leverage these insights to improve the theoretical foundations and practical applications of machine learning in analyzing and predicting dynamical systems.
@@
Most results in AIT, and the original algorithmic probability formulation, are given in terms of binary strings. While AIT is not restricted only to applications involving binary strings, applications involving binary strings are more natural and arguably easier to implement. Hence, in this study, we will digitize trajectories so as to study binary string trajectories (Cf. [45, 27, 43]). We begin by studying simplicity bias in the random logistic map and then investigate the effects of adding measurement noise also. The following section looks at complexity-probability graphs and simplicity bias in the noise-induced regime of the logistic map. Finally, we look at the logistic map in the context of the counter-intuitive claim of induction based on algorithmic probability, that sometimes has more data, all of which follows the same trend, can result in less confidence in our ability to extrapolate the trend.
@@ 2.1.1. AIT and Kolmogorov complexity
At the intersection of computer science and information theory lies algorithmic information theory [21, 22, 23] (AIT). AIT concerns the information content of individual objects, such as binary strings and discrete geometries [40]. The central quantity of AIT is Kolmogorov complexity, $K(x)$, which quantifies the shortest length of a program to produce some object $x$. Alternatively $K(x)$ can be thought of as the size in bits of the compressed version of $x$ (assuming a perfect compressor). An object $x$ which has a complex or random pattern is hard to compress, and so has a high Kolmogorov complexity, while an object $x$ containing simple repeating patterns like $x=01010\dots 10101$ is highly compressible and so has a low Kolmogorov complexity value.
@@ 2.1.2. The coding theorem and algorithmic probability
where $P(x)$ is the probability that an output $x$ is generated by a generic computing device (technically a prefix optimal universal Turing machine) when fed with a random binary program. Thus, high complexity outputs have exponentially low probability, and simple outputs must have high probability. $P(x)$ is also known as the algorithmic probability of $x$ (see Ref. [26] for an overview). AIT results are often difficult to apply directly in real-world contexts, due to a number of issues including the fact that $K(x)$ is formally uncomputable.
@@ 2.1.3. The simplicity bias bound
    The roughly linear upper bound in (ii) follows the simplicity bias bound (see below) with $a\approx 1$
@@
The main motivations for studying simplicity bias are (a) uncovering a ‘universal’ property of input-output maps; (b) identifying a basic and general mechanism for the appearance of simplicity and symmetry in nature; and (c) exploring a fundamentally different way to predict the probability of output shapes as compared to standard methods like frequency sampling. Simplicity bias means that some predictions about output probabilities can be made directly by estimating the complexity of outputs, independently of the underlying system mechanisms and without having to make frequency counts.
@@ 2.2. Digitised map trajectories
AIT results, including algorithmic probability, are typically given in terms of random inputs or ‘programs’ which are computed to produce output strings or patterns. Further, these results of AIT are normally given in terms of binary strings. While dynamical systems may not appear to fit this input-output framework, they can be viewed as input-output functions if the initial value and parameters are considered input ‘programs’ which are then computed to generate ‘output’ dynamical system trajectories. As above, because output pattern complexities and probabilities are more easily calculated if outputs are binary strings, we will digitize the real-valued trajectories into binary strings.
@@
The random logistic map can be viewed as an input-output map because the values of $\mu$ and $x_{0}$ are like ‘inputs’ which are computed (in addition to the noise) to produce a binary sequence trajectory ‘output’. In more detail, the map is first iterated to obtain a sequence of $n$ real values $x_{1},x_{2},x_{3},\dots,x_{n}$ in [0,1]. Similar to the field of symbolic dynamics [54], the real-valued trajectory is digitized to become a binary string output sequence $x$ by applying a threshold, writing 1 if $x_{k}\geq 0.5$ and 0 if $x_{k}<0.5$ [45, 55, 27]. Hence a binary sequence output $x$ of $n$ bits is generated for each input pair $\mu$ and $x_{0}$, and random realization of $\omega_{k}$ for $k=1,2,3,\dots,n$.
@@
Following Ref. [27], we will use $n=25$, but see the same reference for the impact of larger are smaller $n$. In brief, $n$ should be large enough to yield many different patterns and complexity values, but small enough so that decent frequency and probability estimates can be made for up to $2^{n}$ different strings without the need for excessive sampling.
@@ 3.1. Simplicity bias in the random logistic map
In this study, we are primarily interested in values of $\mu$-values for which the trajectories of the deterministic map converge quite quickly to fairly trivial fixed points or periods. If instead $\mu\approx 4.0$ then it is known that there is no simplicity bias in trajectories [27], so adding noise will only serve to increase the complexity and variability of the trajectories, making any kind of bias and therefore simplicity bias, not appear. Note that we use the word ‘bias’ to describe a strongly non-uniform (low entropy) probability distribution, such that certain binary string outputs are much more likely the others.
@@
It is known that if $\mu\in[0.0,1.0]$, then the trajectories of the deterministic logistic map tend to 0 [57]. Because we are adding noise to the trajectories, the trajectories will be prevented from becoming arbitrarily close to zero. If the noise is small enough, then $x_{k}\approx 0<0.5$ so that the resulting digitized binary string (after truncating at 0.5) will be $x=...0000$. If the noise is larger, then it may be that $x_{k}>0.5$ for some $k$, so that more varied patterns including 1s can emerge for $x$. For $\mu\in(1.0,3.0]$ the trajectories of the deterministic logistic map tend to a fixed point with value $1-(1/\mu)$ [58]. If $\epsilon=0.0$, then for any initial value of $x_{0}$ the trajectories will tend to be either $x=\dots 000$ or $x=\dots 111$ depending on $\mu$, which are trivial patterns. Not only are they trivial, they are not varied enough to have different complexity and probability values, and so the corresponding complexity-probability plot is trivial. Therefore, simplicity bias cannot emerge in this scenario, at least if we ignore the initial transient dynamics (but see also Ref. [27] for more on this). However, if we add noise then different and non-trivial patterns can emerge.
@@
Figure 1. Simplicity bias in the digitized random logistic map. Each blue data point corresponds to a different binary digitized trajectory $x$ of length 25 bits. The black line is the upper bound of Eq. (2). (a) $\mu=3.0$ and $\epsilon=0.125$ including transient dynamics; (b) $\mu=1.0$ and $\epsilon=0.375$ including transient dynamics; and (c) $\mu=1.0$ and $\epsilon=0.375$ excluding transient dynamics.
@@
The parameter values used in the numerical experiments are $\mu=$ 0.0, 1.0, 2.0, and 3.0 in combination with ϵ=italic-$\epsilon=$ 0.125, 0.25, 0.375, and 0.5. For many of these parameter combinations, we can expect no simplicity bias in the binary trajectories, due to the following: If the noise $\epsilon$ is large, then the trajectory is essentially a random erratic pattern across the interval [0,1] with little correlation between adjacent bits, yielding a roughly uniform distribution. Therefore, there is unlikely to be bias for some specific patterns, and consequently no bias for simple patterns. On the other hand, if $\epsilon$ is very small, then the ‘kicks’ imposed by the noise will not be large enough to prevent the trajectories from converging. As examples, if $\mu=1.5$ and $\epsilon\approx 0$, then the trajectory values $x_{k}$ will fluctuate around the value $0.51-(1/\mu)=0.33<$ 0.5, yielding a binary trajectory $x=...0000$. If $\mu=2.0$, then the fixed point would be $0.51-(1/\mu)=$ 0.5, so even small fluctuations around this value would yield almost random binary strings $x$. From this reasoning, we can conclude that to observe simplicity bias in the random logistic map requires having a balance of noise that is not too large as to overly randomize the trajectories, and not too small as to allow the trajectories to converge to a fixed point with only modest fluctuations around $1-(1/\mu)$.
@@
Figure 1 shows some examples of simplicity bias. The different binary string outputs that appeared in sampling are shown as blue dots. The black line bound of Eq. (2) is also depicted. Figure 1(a) shows $\mu=3.0$ and $\epsilon=0.125$, and the initial dynamics are included. There is simplicity bias because we see that some low complexity patterns have high probability. Further, we see a roughly linear upper bound decay in the probabilities, as complexity $\tilde{K}(x)$ increases. However, the gradient appears to be less steep than the $a=1.0$ default prediction suggests. The reason that the slope is not accurate may be due to undersampling because estimating the slope $a$ assumes that nearly all of the possible strings have appeared in the sample, which may be overly coarse. Another reason may be due to the introduction of random noise (see next section for more on this). Despite these considerations, it is still clear that simplicity bias is observed. Without the additive noise, the trajectory would have converged to $0.671-1/\mu=$ 0.67 and shown no simplicity bias.
In Figure 1(b) we see data for $\mu=1.0$ and $\epsilon=0.375$, and the initial dynamics are included. Again there is a clear simplicity bias. Without the additive noise, the trajectory would have converged to $0.01-1/\mu=$ 0.0 and shown no simplicity bias. In addition to the inverse relation between probability and complexity, even the gradient of the decay of the upper bound is not strongly divergent from the black line default prediction. In Figure 1(c) we see data again for $\mu=1.0$ and $\epsilon=0.375$, but this time after ignoring the first 50 iterations (to exclude the transient dynamics). Here we see simplicity bias in the form of an inverse relation between probability and complexity. Having said that, the simplicity bias in panel (c) is not as pronounced as in panel (b).

### 08 Apr 2024 v2

@@ Abstract.
Simplicity bias is an intriguing phenomenon prevalent in various input-output maps, characterized by a preference for simpler, more regular, or symmetric outputs. Notably, these maps typically feature high-probability outputs with simple patterns, whereas complex patterns are exponentially less probable. This bias has been extensively examined and attributed to principles derived from algorithmic information theory and algorithmic probability. In a significant advancement, it has been demonstrated that the renowned logistic map $x_{k+1}=\mu x_{k}(1-x_{k})$, a staple in dynamical systems theory, and other one-dimensional maps exhibit simplicity bias when conceptualized as input-output systems. Building upon this work, our research delves into the manifestations of simplicity bias within the random logistic map, specifically focusing on scenarios involving additive noise. This investigation is driven by the overarching goal of formulating a comprehensive theory for the prediction and analysis of time series.
Our primary contributions are multifaceted. We discover that simplicity bias is observable in the random logistic map for specific ranges of $\mu$ and noise magnitudes. Additionally, we find that this bias persists even with the introduction of small measurement noise, though it diminishes as noise levels increase. Our studies also revisit the phenomenon of noise-induced chaos, particularly when $\mu=3.83$, revealing its characteristics through complexity-probability plots. Intriguingly, we employ the logistic map to illustrate a paradoxical aspect of data analysis: more data adhering to a consistent trend can occasionally lead to reduced confidence in extrapolation predictions, challenging conventional wisdom.
@@ 1. Introduction
Dynamical systems appear in many fields including financial, engineering, meteorological, medical, and other areas. Because of this, developing methods to predict or understand the behavior and trajectories of dynamical systems is an important area of applied mathematics research. Historically, such predictions have been based on modeling the relevant system (e.g., with ODEs). More recently, however, machine learning advances have given tools for the prediction and analysis of dynamical systems. The machine learning approach instead relies on directly extracting patterns from data rather than detailed modeling of the underlying mechanics [1, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21].
@@
Numerous studies of simplicity bias have been made in input-output maps, in which a general inverse relationship between the complexity of outputs and their respective probabilities has been observed [33, 34]. This work is based on information theory and computation results and was derived from arguments inspired by AIT, and can be viewed as an attempt to apply algorithmic probability in real-world settings [33]. The simplicity bias upper bound [33] gives a way to bound the probability $P(x)$ of observing output pattern $x$ based essentially only on estimating the Kolmogorov complexity of the pattern in $x$. Simplicity bias analysis has been applied in many contexts, including machine learning and deep neural networks [35, 36, 37, 38, 39, 40, 41]. Simplicity bias and algorithmic probability are closely related to Occam’s razor [42], a fundamental basis of scientific reasoning and model selection, that simpler theories/models should be preferred over more complex theories/models provided they explain the data equally well.
The goal of this paper is to bridge the gap between machine learning and algorithmic information theory using dynamical systems, particularly focusing on the concept of simplicity bias within random logistic maps. We aim to extend the understanding of simplicity bias to encompass the inherently stochastic nature of real-world systems as modeled by random dynamical systems. Our approach is twofold: first, to elucidate the manifestation of simplicity bias in the random logistic map under various conditions and noise levels, thereby providing insights into the behavior of these systems; second, to leverage these insights to improve the theoretical foundations and practical applications of machine learning in analyzing and predicting dynamical systems.
@@
Most results in AIT, and the original algorithmic probability formulation, are given in terms of binary strings. While AIT is not restricted only to applications involving binary strings, applications involving binary strings are in a sense more natural and arguably easier to implement. Hence, in this study, we will digitize trajectories so as to study binary string trajectories (Cf. [47, 29, 45]). We begin by studying simplicity bias in the random logistic map and then investigate the effects of adding measurement noise also. The following section looks at complexity-probability graphs and simplicity bias in the noise-induced regime of the logistic map. Finally, we look at the logistic map in the context of the counter-intuitive claim of induction based on algorithmic probability, that sometimes more data, all of which follows the same trend, can result in less confidence in our ability to extrapolate the trend.
@@ 2.1.1. AIT and Kolmogorov complexity
At the intersection of computer science and information theory lies algorithmic information theory [23, 24, 25] (AIT). AIT concerns the information content of individual objects, such as binary strings, integers, and discrete geometries [42]. The central quantity of AIT is Kolmogorov complexity, $K(x)$, which quantifies the shortest length of a program to produce some object $x$. Alternatively $K(x)$ can be thought of as the size in bits of the compressed version of $x$ (assuming a perfect compressor). An object $x$ which has a complex or random pattern is hard to compress, and so has a high Kolmogorov complexity, while an object $x$ containing simple repeating patterns like $x=01010\dots 10101$ is highly compressible and so has a low Kolmogorov complexity value.
@@ 2.1.2. The coding theorem and algorithmic probability
where $P(x)$ is the probability that an output $x$ is generated by a generic computing device (technically a prefix optimal universal Turing machine) when fed with a random binary program. Thus, high complexity outputs have exponentially low probability, and simple outputs with small $K(x)$ values must have high probability. $P(x)$ is also known as the algorithmic probability of $x$ (see Ref. [28] for an overview). AIT results are often difficult to apply directly in real-world contexts, due to a number of issues including the fact that $K(x)$ is formally uncomputable.
@@ 2.1.3. The simplicity bias bound
    A complexity-probability plot as in (ii) but with the linear upper bound following the simplicity bias bound Eq. (2) (see below) with $a\approx 1$
@@
The main motivations for studying simplicity bias are (a) uncovering a ‘universal’ property of input-output maps; (b) identifying a basic and general mechanism for the appearance of simplicity and symmetry in nature; and (c) exploring a fundamentally different way to predict the probability of output shapes as compared to standard methods like frequency sampling. Simplicity bias means that some predictions about output probabilities can be made directly by estimating the complexity of outputs, largely independently of the underlying system mechanisms and without having to make frequency counts.
@@ 2.2. Digitised map trajectories
AIT results, including algorithmic probability, are typically given in terms of random inputs or ‘programs’ which are computed to produce output strings or patterns. Further, these results of AIT are normally given in terms of binary strings. While dynamical systems may not appear to fit this input-output framework, they can be viewed as input-output functions if the initial value and parameters are considered input ‘programs’ which are then computed to generate ‘output’ dynamical system trajectories. As above, because output pattern complexities and probabilities are more straightforwardly calculated if outputs are binary strings, we will digitize the real-valued trajectories into binary strings.
@@
The random logistic map can be viewed as an input-output map because the values of $\mu$ and $x_{0}$ are like ‘inputs’ which are computed (in addition to the noise) to produce a binary sequence trajectory ‘outputs’. In more detail, the map is first iterated to obtain a sequence of $n$ real values $x_{1},x_{2},x_{3},\dots,x_{n}$ in [0,1]. Similar to the field of symbolic dynamics [56], the real-valued trajectory is digitized to become a binary string output sequence $x$ by applying a threshold, writing 1 if $x_{k}\geq 0.5$ and 0 if $x_{k}<0.5$ [47, 57, 29]. Hence a binary sequence output $x$ of $n$ bits is generated for each input pair $\mu$ and $x_{0}$, and random realization of $\omega_{k}$ for $k=1,2,3,\dots,n$.
@@
Following Ref. [29], we will use $n=25$, but see the same reference for the impact of larger or smaller $n$. In brief, $n$ should be large enough to yield many different patterns and complexity values, but small enough so that decent frequency and probability estimates can be made for up to $2^{n}$ different strings without the need for excessive sampling.
@@ 3.1. Simplicity bias in the random logistic map
In this study, we are primarily interested in $\mu$-values for which the trajectories of the deterministic map converge quite quickly to fairly trivial fixed points or periods. If instead $\mu\approx 4.0$ then it is known that there is no simplicity bias in trajectories [29], so adding noise will only serve to increase the complexity and variability of the trajectories, making any kind of bias and therefore simplicity bias, not appear. Note that we use the word ‘bias’ to describe a strongly non-uniform (low entropy) probability distribution, such that certain binary string outputs are much more likely the others.
@@
It is known that if $\mu\in[0.0,1.0]$, then the trajectories of the deterministic logistic map tend to 0 [59]. Because we are adding noise to the trajectories, the trajectories will be prevented from becoming arbitrarily close to zero. If the noise is small enough, then $x_{k}\approx 0<0.5$ so that the resulting digitized binary string (after truncating at 0.5) will be $x=...0000$. If the noise is larger, then it may be that $x_{k}>0.5$ for some $k$, so that more varied patterns including 1s can emerge. For $\mu\in(1.0,3.0]$ the trajectories of the deterministic logistic map tend to a fixed point with value $1-(1/\mu)$ [60]. If $\epsilon=0.0$, then for any initial value of $x_{0}$ the trajectories will tend to be either $x=\dots 000$ or $x=\dots 111$ depending on $\mu$, which are trivial patterns. Not only are they trivial, they are not varied enough to have different complexity and probability values, and so the corresponding complexity-probability plot is trivial. Therefore, simplicity bias cannot emerge in this scenario, at least if we ignore the initial transient dynamics (but see also Ref. [29] for more on this). However, if we add noise then different and non-trivial patterns can emerge.
@@
Figure 1. Simplicity bias in the digitized random logistic map. Each blue data point corresponds to a different binary digitized trajectory $x$ of length 25 bits. The black line is the upper bound of Eq. (2) with $a=1$ as a default prediction. (a) $\mu=3.0$ and $\epsilon=0.125$ including transient dynamics; (b) $\mu=1.0$ and $\epsilon=0.375$ including transient dynamics; and (c) $\mu=1.0$ and $\epsilon=0.375$ excluding transient dynamics.
@@
The parameter values used in the numerical experiments are $\mu=$ 0.0, 1.0, 2.0, and 3.0 in combination with ϵ=italic-$\epsilon=$ 0.125, 0.25, 0.375, and 0.5. For many of these parameter combinations, we can expect no simplicity bias in the binary trajectories, due to the following: If the noise $\epsilon$ is large, then the trajectory is essentially a random erratic pattern across the interval [0,1] with little correlation between adjacent bits, yielding a roughly uniform distribution over strings. Therefore, there is unlikely to be bias for some specific patterns, and consequently no bias for simple patterns. On the other hand, if $\epsilon$ is very small, then the ‘kicks’ imposed by the noise will not be large enough to prevent the trajectories from converging. As examples, if $\mu=1.5$ and $\epsilon\approx 0$, then the trajectory values $x_{k}$ will fluctuate around the value $0.51-(1/\mu)=0.33<$ 0.5, yielding a binary trajectory $x=...0000$. If $\mu=2.0$, then the fixed point would be $0.51-(1/\mu)=$ 0.5, so even small fluctuations around this value would yield almost random binary strings $x$. From this reasoning, we can conclude that to observe simplicity bias in the random logistic map requires having a balance of noise that is not too large as to overly randomize the trajectories, and not too small as to allow the trajectories to converge to a fixed point with only modest fluctuations around $1-(1/\mu)$.
@@
Figure 1 shows some examples of simplicity bias. The different binary string outputs that appeared in sampling are shown as blue dots. The black line bound of Eq. (2) is also depicted with $a=1$ (the default prediction). Figure 1(a) shows $\mu=3.0$ and $\epsilon=0.125$, and the initial dynamics are included. There is simplicity bias because we see that some low complexity patterns have high probability. Further, we see a roughly linear upper bound decay in the probabilities, as complexity $\tilde{K}(x)$ increases. However, the gradient appears to be less steep than the $a=1.0$ default prediction suggests. The slope of the upper bound is -0.27. The reason that the slope is not accurate may be due to undersampling because estimating the slope $a$ assumes that nearly all of the possible strings have appeared in the sample, which may be overly coarse. Another reason may be due to the introduction of random noise (see next section for more on this). Despite these considerations, it is still clear that simplicity bias is observed. Without the additive noise, the trajectory would have converged to $0.671-1/\mu=$ 0.67 and shown no simplicity bias.
In Figure 1(b) we see data for $\mu=1.0$ and $\epsilon=0.375$, and the initial dynamics are included. Again there is clear simplicity bias. Without the additive noise, the trajectory would have converged to $0.01-1/\mu=$ 0.0 and shown no simplicity bias. In addition to the inverse relation between probability and complexity, even the gradient of the decay of the upper bound is not strongly divergent from the black line default prediction. The slope of the upper bound is -0.32. In Figure 1(c) we see data again for $\mu=1.0$ and $\epsilon=0.375$, but this time after ignoring the first 50 iterations (to exclude the transient dynamics). Here we see simplicity bias in the form of an inverse relation between probability and complexity. Having said that, the simplicity bias in panel (c) is not as pronounced as in panel (b). The slope of the upper bound is -0.37.